OpenStudy (anonymous):

Find the limit:

OpenStudy (anonymous):

\[(\int\limits_{0}^{x}(3t-1)^{10} dt)/3x\]

OpenStudy (anonymous):

as x approaches 0 I'm confused because the question has two variables: x and t. How should I approach this?

OpenStudy (watchmath):

0

OpenStudy (mr.math):

Evaluate the limit in the top first.

OpenStudy (watchmath):

sorry, I didn't see 3x :)

OpenStudy (watchmath):

Use L'Hospital rule

OpenStudy (mr.math):

Oh right. L'Hopital's rule is the best choice: \[\frac{d}{dx}\int\limits_0^x(3t-1)^{10}dt=(3x-1)^{10}.\]

OpenStudy (mr.math):

The derivative of the bottom is obviously 3. Hence the limit is \(\frac{3(0)-1)^{10}}{3}=\frac{1}{3}\).

OpenStudy (anonymous):

This makes sense, but how did you know to use l'hopital's rule?

OpenStudy (mr.math):

If we plug x=0, you will get \(\frac{0}{0}\). If you plug x in the top, you get an integral from 0 to 0, which results in a value of 0. I think it's very obvious in the denominator.

OpenStudy (anonymous):

oh okay, thanks a ton!